OpenStudy (anonymous):

If a, b, c are in A.P show that 1/ab, 1/bc, 1/ca are also in A.P.

OpenStudy (anonymous):

since a,b,c are in AP let d be the common difference between them, so we hav b-a=d c-b=d 1/bc-1/ab=1/b(1/c-1/a)=(a-c)/abc=-2d/abc 1/ca-1/bc=(b-a)/abc=d/abc which shows theyr not in AP !